OpenStudy (anonymous):

x + y = 6 2x + y = 4 Solve the system of equations by substitution. (-2, 8) (2, -8) no solution coincident

OpenStudy (nurali):

x + y = 6----(1) 2x + y = 4---(2) in equation (1) x+y=6 x=6-y put in equation (2) 2(6-y)+y=4 12-2y+y=4 12-y-4=0 y=8 put these value of y in equation (1) x+8=6 x=6-8 x=-2 so solution (x,y)=(-2,8)

OpenStudy (anonymous):

what about this one? 7x - 2 = 2y 3x = 2y - 1 Solve the system of equations by substitution. (3/4, 13/8) (1/4, -1/8) no solution coincident

OpenStudy (nurali):

7x-2=2y---(1) 3x+1=2y---(2) SUBTRACT EQUATION (1) and (2) ---------------- 4X-3=0 4X=3 X=3/4 put the value of x in equation (1) 7*3/4-2=2Y 21/4-2=2Y 2Y=21/4-8/4 2Y=13/4 Y=(13/4)/2 Y=13/4*1/2 Y=13/8 so solution (x,y)=(3/4,13/8)

OpenStudy (anonymous):

ok i got one last one. 8x = 2y + 5 3x = y + 7 Solve the system of equations by substitution. (-9/2, -41/2) (-1, -9/2) no solution coincident

OpenStudy (nurali):

8x = 2y + 5----(1) 3x = y + 7----(2) in equation (2) Y=3X-7 put in equation (1) 8X=2(3X-7)+5 8X=6X-14+5 8X-6X=-9 2X=-9 X=-9/2 X=-4.5 put the value of x in equation (1) 8*-4.5=2Y+5 -36=2Y+5 2Y=-36-5 2Y=-41 Y=-41/2 Y=-20.5 so solution (x,y)=(-4.5,-20.5)

OpenStudy (anonymous):

ok but the last answer does not match any of my options here they are. a.(-9/2, -41/2) b.(-1, -9/2) c.no solution d.coincident

OpenStudy (nurali):

no solution